OpenStudy (anonymous):

will give medal and fan Simplify 5 • 2 • (–4) + 3 • 6 (Points : 2) –258 –22 4 58

Directrix (directrix):

5 • 2 • (–4) + 3 • 6 This can be grouped for easier understanding. [ 5 • 2 • (–4) ] + [ 3 • 6 ]

Directrix (directrix):

@skoob1 What do you get when you multiply 5 times 2 times -4

Directrix (directrix):

Post what you get. It is okay if it is wrong. That is what the delete button is for.

OpenStudy (anonymous):

-40

Directrix (directrix):

Right. What do you get when you multiply 3 • 6

OpenStudy (anonymous):

18

Directrix (directrix):

Okay, one more step.

Directrix (directrix):

Here is the whole problem put back together: 5 • 2 • (–4) + 3 • 6 = [ 5 • 2 • (–4) ] + [ 3 • 6 ] = - 40 + 18 =

OpenStudy (anonymous):

-22, wasnt so hard after all

Directrix (directrix):

Correct. It's often easier to separate a problem into parts. You did a good job.